Summary for week of July 30 - Aug 5

Julie Whitmore (jaws@fnpxsr)
Fri, 13 Aug 1999 16:35:17 CDT

Summary of E832 Week 8: July 30-Aug 5 Runs: 14061-14122

Hours of beam used/delivered : 101.19/124.05 = 0.82

Vac00 (week) : 325K
34K @ 2E12 (includes 15-20M offline vac Ke3 w/ PS=1)
223K @ 4E12
68K @ 7E12
Vac00 (99run) : 2.22 million

Comments:
Most beam loss due to ComEd shutdown and feeder 37 trip.

KTEV used only 82% of the available beam (worst week for KTeV99)
KTEV loss due mainly to LV power supply problems (CsI & DC-OR),
recover from feeder 37 trip (SWIC and absorber problems),
HCC, and DAQ related problems (drt_fresh_start failure due to
disks full/kumquat reset/DAMP probs; bad tape in ktev1/drv2;
crate 53 problem, ktev4 DDD problem).

hours of beam:
Shift used/delivered comment
---------------------------------------------------------------------
Fri July 30 OWL: 6.5/7.0
DAY: 1.0/1.0 ComEd shutdown
EVE: 0.0/0.0 Begin recovery from shutdown

Sat July 31 OWL: 2.25/3.0 TeV quench problem. CsI LV sector 8 trip.
DAY: 5.25/8.25 CsI LV sector 8 trip.
EVE: 5.0/5.0 TeV TECAR problems.

Sun Aug 1 OWL: 5.7/5.7 TeV problems; bad beam structure
DAY: 8.25/8.5 SEM ~3.8E12
EVE: 5.0/8.5 Swap DC-OR power supply

Mon Aug 2 OWL: 5.8/6.0 SEM ~ 3.9E12
DAY: 4.6/4.6 TeV problems; Begin 2E12 Ke3hog running
EVE: 7.8/8.5 2E12 Ke3hog

Tue Aug 3 OWL: 5.5/5.5 2E12 Ke3hog; Linac problems
DAY: 7.25/7.75 End 2E12 Ke3hog running; back to 4E12
EVE: 7.3/8.5 feeder 37 trip

Wed Aug 4 OWL: 0.0/0.0 Trip recovery;SWIC/absorber prob;3.5h bad beam
DAY: 7.0/8.0 Begin 6.8E12 running
EVE: 5.54/7.75

Thu Aug 5 OWL: 3.2/4.0 Continue 6.8E12 (Ke3 PS:30-->50)
DAY: 4.75/8.5 drt_fresh_start problems;bad tape
EVE: 3.5/8.0 HCC, crate 53, and DDD4 problems
---------------------------------------------------------------------
TOTAL 101.19/124.05 = 0.82
---------------------------------------------------------------------